The floor of a shed has an area of 54 square feet. Find the length and width if the length is 3 feet less than twice its width.
Answer by checkley71(8403)

AREA=LENGTH*WIDTH
54=(2W-3)W
54=2W^2-3W
2W^2-3W-54=0
(2W+9)(W-6)=0
2W+9=0
2W=-9
W=9/-2
W=-4.5 NON SOLUTION
W-6=0
W=6 SOLUTION
L=2*6-3
L=12-3
L=9 SOLUTION
PROOF
54=6*9
54=54